Required information

Problem 2-4 (Algo) Accounting cycle; adjusting entries through post-closing trial balance [LO2-4, 2-6, 2-7, 2-8]

[The following information applies to the questions displayed below.]


Pastina Company sells various types of pasta to grocery chains as private label brands. The company's reporting year-end is December 31. The unadjusted trial balance as of December 31, 2021, appears below.

   

Account Title Debits Credits
Cash 34,100
Accounts receivable 42,000
Supplies 2,500
Inventory 62,000
Notes receivable 22,000
Interest receivable 0
Prepaid rent 1,900
Prepaid insurance 8,000
Office equipment 88,000
Accumulated depreciation 33,000
Accounts payable 33,000
Salaries payable 0
Notes payable 52,000
Interest payable 0
Deferred sales revenue 3,000
Common stock 74,000
Retained earnings 33,500
Dividends 6,000
Sales revenue 156,000
Interest revenue 0
Cost of goods sold 80,000
Salaries expense 19,900
Rent expense 12,000
Depreciation expense 0
Interest expense 0
Supplies expense 2,100
Insurance expense 0
Advertising expense 4,000
Totals 384,500 384,500

Information necessary to prepare the year-end adjusting entries appears below.

  1. Depreciation on the office equipment for the year is $11,000.
  2. Employee salaries are paid twice a month, on the 22nd for salaries earned from the 1st through the 15th, and on the 7th of the following month for salaries earned from the 16th through the end of the month. Salaries earned from December 16 through December 31, 2021, were $1,250.
  3. On October 1, 2021, Pastina borrowed $52,000 from a local bank and signed a note. The note requires interest to be paid annually on September 30 at 12%. The principal is due in 10 years.
  4. On March 1, 2021, the company lent a supplier $22,000 and a note was signed requiring principal and interest at 8% to be paid on February 28, 2022.
  5. On April 1, 2021, the company paid an insurance company $8,000 for a one-year fire insurance policy. The entire $8,000 was debited to prepaid insurance.
  6. $770 of supplies remained on hand at December 31, 2021.
  7. A customer paid Pastina $3,000 in December for 1,250 pounds of spaghetti to be delivered in January 2022. Pastina credited deferred sales revenue.
  8. On December 1, 2021, $1,900 rent was paid to the owner of the building. The payment represented rent for December 2021 and January 2022 at $950 per month. The entire amount was debited to prepaid rent.

Problem 2-4 (Algo) Part 4

4. Prepare an income statement and a statement of shareholders’ equity for the year ended December 31, 2021, and a classified balance sheet as of December 31, 2021. Assume that no common stock was issued during the year and that $6,000 in cash dividends were paid to shareholders during the year.
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Due to character limit I am attaching the image for adjusted Trial Balance.
Pastina
Working for Adjusting Entry Amount $ Note
Note payable           52,000.00 A
Interest rate 12% B
Annual Interest amount             6,240.00 C=A*B
Monthly Interest                 520.00 D=C/12
Interest from Oct to Dec (3 months)             1,560.00 E=D*3
Note receivable           22,000.00 F
Interest rate 8% G
Annual Interest amount             1,760.00 H=F*G
Monthly Interest                 146.67 I=H/12
Interest from March to Dec (10 months)             1,466.67 J=I*10
Prepaid Insurance             8,000.00 K
For the period (months)                   12.00 L
Monthly Insurance                 666.67 M=K/L
Insurance from Apr to Dec (9 months)             6,000.00 N=M*9
Supplies Expense
Supplies - Opening Balance             2,500.00
Less: Supplies - Closing Balance                 770.00
Supplies Expense             1,730.00 O
Prepaid Rent             1,900.00 P
Rent for December                 950.00 Q
Adjusting Entries
Account Debit $ Credit $
Depreciation Expense        11,000.00
Accumulated Depreciation- Equipment        11,000.00
Salaries Expense          1,250.00
Salaries Payable          1,250.00
Interest Expense          1,560.00
Interest Payable          1,560.00
Interest Receivable          1,466.67
Interest Revenue          1,466.67
Insurance Expense          6,000.00
Prepaid Insurance          6,000.00
Supplies Expense          1,730.00
Supplies          1,730.00
Rent Expense             950.00
Prepaid Rent             950.00

Unadjusted Adjustments Adjusted Trial Balance Debit Credit Debit Credit Debit Credit Cash 34,100.00 34,100.00 Accounts Receiv

Income Statement
Particulars Amount $ Amount $
Sales Revenue         156,000.00
Less: returns and allowances                          -  
Net Sales         156,000.00
Cost of goods sold           80,000.00
Gross Profit           76,000.00
Salaries Expense           21,150.00
Rent Expense           12,950.00
Depreciation Expense           11,000.00
Supplies Expense             3,830.00
Insurance Expense             6,000.00
Advertising Expense             4,000.00
Operating Expense           58,930.00
Operating Income           17,070.00
Interest Income             1,466.67
Interest expense            (1,560.00)
Other expense and income                 (93.33)
Net Income           16,976.67
Statement of shareholder's equity Common Stock Retained Earnings Total
Balance at January 1, 2021           74,000.00           33,500.00 107,500.00
Net Income for the year                          -             16,976.67     16,976.67
Dividend                          -               6,000.00       6,000.00
Balance at December 31, 2021           74,000.00           44,476.67 118,476.67
Balance Sheet Amount $ Amount $
Current Assets
Cash           34,100.00
Accounts Receivable           42,000.00
Inventory           62,000.00
Supplies                 770.00
Notes Receivable           22,000.00
Interest Receivable             1,466.67
Prepaid Rent                 950.00
Prepaid Insurance             2,000.00
Total Current Assets         165,286.67
Non Current Assets
Office Equipment